OpenStudy (tatumlee):

Solve the system of equations by substitution x=-4y x+5y=2 a.(-8,2) b.(8,-2) c.(2/9,-8/9) d.(-12,3) How do I do this problem?

pooja195 (pooja195):

It tells you the vale of x right ? So just plug in the value of x into equation 2 -4y+5y=2 Now solve the equation for y tell me what you get

pooja195 (pooja195):

Ye solve for y -4+5=?

OpenStudy (tatumlee):

1

pooja195 (pooja195):

Ok so 1y=2 Now divide both sides by 1 2/1=?

OpenStudy (tatumlee):

it would just be 2 i think?

pooja195 (pooja195):

Good so y=2 Now to solve for x... Now that we know the value of y we can plug in 2 (into equation 1) to figure out the x value -4 x 2 = ?

OpenStudy (tatumlee):

so -8?

pooja195 (pooja195):

Yes so x=-8 y=2 which coordinate matches the answer?

OpenStudy (tatumlee):

so the answer is a? and what do you mean?

pooja195 (pooja195):

A is correct :)

OpenStudy (tatumlee):

Thanks again!

pooja195 (pooja195):

:)